0.01 mole of iodoform reacts with excess Ag powder to produce a gas whose volume at NTP is

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Hiii satyam 
chemical reaction of above question is :
2CHI3 + 6Ag = C2H2 + 6AgI
0.01 moe of iodoform forms 0.005 mole of acetylene gas.
n= 0.005 , R=0.082 atm.litre/J.mol.K 
T= 298 K , P= 1 atm
PV = nRT
1*V = 0.005*0.0082*298
V= 0.122litre or 122 ml ans